OpenStudy (anonymous):

what is the equation of the line in slope-intercept form that passes throught the point 1 3 and is perpendicular to the line y=-x+2 a) y=x-4 b) y=x+2 c) y=-x+2 d) y=-x-4 HELP !!!!!!

OpenStudy (anonymous):

let the line is y=x+k put 1,3 3=1+k k=2 y=x+2